The Creighton family owns a vacation rental which generates $10530 in income each quarter. The family outsources maintenance to another company which charges $1065 per quarter, with the amount increasing by $26 each subsequent quarter. What is the equivalent uniform quarterly worth of the vacation rental when looked at over a 17-year planning period? Use a nominal annual interest rate of 8%, compounded quarterly to make the calculation.
Answer #1

Rental income every quarter = 10530

T=17 yrs = 17*4 = 68 months

i = 8% compounded quarterly

i = 2% per quarter

Maintenance = 1065 in the first quarter and increasing by 26 each quarter.

we will assume a uniform series of 1065 from 1st to 68th quarter and gradient series for an increase in maintenance amount of 26

We will use factor (A/G, 2%,68)

uniform rental worth = 10530-1065-26* (A/G, 2%,68)

= 9465 - 26*26.09211

= 8786.60